# SGSST: Scaling Gaussian Splatting Style Transfer

### v1

@@ 1 Introduction
In short, our approach is the first method that allows UHR style transfer directly to 3DGS. It produces high visual quality results and relies on optimizing a single multiscale loss. The simplicity of our approach ensures its reproducibility (source code is provided as supplementary material). Being optimization-based, SGSST’s main limitation is a fairly large training time that is two to eight times longer than the initial 3DGS training depending on the image resolution. Yet, considering the high quality of the results and their reproducibility, this contribution is valuable for AR/VR applications that require high visual quality for their user experience.
@@ 2 Related work
A few recent contributions show that 3DGS is a promising framework for 3D scene style transfer. Saroha et al [33] propose a solution for universal style transfer of a given Gaussian splatting 3D scene. The method processes the colors of Gaussians with a tiny MLP trained using fast AdaIN [15] and relying on a multi-resolution hash grid representation [29]. StyleGaussian [26] is a concurrent approach that relies on transferring encoding of VGG features to each Gaussian and applying AdaIN to these features. The new Gaussian features are then decoded into an RGB color by processing the K-nearest neighbors of the Gaussians. After training, both methods allow for instantaneous stylization with any style image, but the visual quality of the results is quite low, since high-resolution details such as brushstrokes patterns are not transferred.
StylizedGS [46] is an optimization based method that extends ARF [48] to 3DGS. It uses a training loss made of six terms combined with a color transfer preprocessing, the loss being designed to minimize changes in the Gaussian splatting geometry while letting the style evolve via VGG NN matching. $\mathcal{G}$-Style [22] also is an optimization based approach that further uses a CLIP-based loss [31]. While these approaches produce slightly better results than ARF [48], both methods are limited in content and style resolution due to the use of nearest-neighbor matching of VGG features.
None of the current state of the art deals with HR style transfer. The stylization is either fast but very approximate due to AdaIn [33, 26] or unable to use HR style images due to NNFM [48, 46, 22]. To the best of our knowledge SGSST is the first procedure allowing high quality transfer for 3DGS that is trained and rendered at UHR.
@@ 3.1 Gaussian splatting representation
Starting from a multiview training set of images $\{u_{i}\}_{i=1}^{N_{\mathrm{views}}}$ of a static scene accompanied with corresponding calibrated cameras $\{\mathcal{C}_{i}\}_{i=1}^{N_{\mathrm{views}}}$ computed by structure from motion [34], the 3DGS algorithm [20] trains a set of colored 3D Gaussians $\{\mathcal{G}_{j}\}_{j=1}^{N_{\mathrm{Gaussians}}}$ so that they represent the 3D scene from any camera position. Each Gaussian $\mathcal{G}_{j}$ is represented by a finite set of features: a center position $\mu_{j}$, a covariance matrix $\Sigma_{j}$ (encoded by a scaling diagonal matrix and a rotation matrix), an opacity $\alpha_{j}$ and a view dependent color function $c_{j}$. These parameters are used in a volumetric renderer that determines the color by summing the contribution of each Gaussian that intersects a ray with direction $(\theta,\varphi)$ via $\alpha$-blending (see e.g. [28]). The contribution of a Gaussian is its color $c_{j}(\theta,\varphi)$ multiplied by an opacity $\sigma_{j}$ defined as the maximal opacity $\alpha_{j}$ times the unnormalized Gaussian density at the ray position [20]. Thus, the resulting color $C$ for the ray is
@@ 3.2 Style transfer loss for UHR images
Figure 2: Overview of SGSST. Starting from a pretrained realistic 3DGS scene [20], we optimize the colors of each Gaussian using the new multiscale SOS loss (involving $n_{\mathrm{s}}=3$ scales in the illustration). Computing the gradient wrt the loss is feasible for UHR images thanks to the SPST partition-based implementation [10]. Multiscale gradient stacking is used at the node of the rendered image to perform only one backpropagation per iteration through the 3DGS rendering pipeline.
@@
where $L_{\mathrm{content}}(x;u)=\lambda_{\mathrm{c}}\left\|V^{L_{\mathrm{c}}}(x)-V^{L_{\mathrm{c}}}(u)\right\|^{2}$, with $\lambda_{\mathrm{c}}>0$ and the style loss is defined by
@@
where $E_{\mathrm{style}}^{L}(x;v)$ is a linear combination of the mean square error between the VGG19 statistics of $V^{L}(x)$ and the one of the style features $V^{L}(v)$ [10]. While only the Gram matrices were originally used [11], it has been shown that adding control for the mean and standard deviation corrects some style transfer color artefacts [10] previously identified in the litterature  [35, 32, 14].
To obtain high-quality style transfer for HR images one needs to optimize the style transfer loss using several scales and a coarse-to-fine approach [12]. Indeed, if one applies style transfer on the highest resolution only, the changes within the content image are limited to local texture and the results does not convey a painting aspect. Due to the use of VGG19 features, computing the gradient of $E_{\mathrm{transfer}}(x;u,v)$ with respect to (wrt) $x$ is a priori memory prohibiting for UHR images, however SPST [10] provides a constant memory algorithm for evaluating this gradient via a grid partition.
@@ 4.2 Multiscale Style Transfer Loss
In the end, the stylization of UHR 3DGS is conducted by solving for
@@ 4.3 Implementation details
The 2D SPST method [10] provides the gradient of each term $E_{\mathrm{transfer}}(x^{\downarrow 2^{s}};u^{\downarrow 2^{s}},v^{\downarrow 2^{s}})$ wrt to the input $x^{\downarrow 2^{s}}=\mathcal{R}(\mathcal{C}_{i};\Theta)^{\downarrow 2^{s}}$. We first backpropagate each of these gradients through the downscaling operator and stack them at the level of the rendered image $\mathcal{R}(\mathcal{C}_{i};\Theta)$. When each of the $n_{\mathrm{s}}$ scales has been treated, the stacked gradient is equal to the gradient of the full loss $E_{\mathrm{SOS}}(\mathcal{R}(\mathcal{C}_{i};\Theta);u_{i},v)$ wrt to the rendered image $\mathcal{R}(\mathcal{C}_{i};\Theta)$. This gradient is then backpropagated through the Gaussian rendering pipeline to obtain the gradient wrt the Gaussian colors $\Theta_{\mathrm{color}}$ (see both orange arrows in Figure 2). Using this strategy we only backpropagate one time per iteration through the 3DGS rendering pipeline instead of $n_{\mathrm{s}}$ times.
@@
The number of scales is chosen between $1$ and $4$ depending on the size of the original images (we always consider images with sides larger than 224 for VGG19 statistics to be reliable).
@@
All our experiments have been conducted using the same SOS loss and training procedure, making our approach parameterless and fully reproducible. Our implementation is based on the public source codes [^1] https://github.com/graphdeco-inria/gaussian-splatting; https://github.com/bgalerne/scaling_painting_style_transfer for 3DGS [20] training and SPST [10]. Our PyTorch source code is available at https://github.com/JianlingWANG2021/SGSST.
@@ 5 Experiments
Figure 4: Comparison of SGSST (ours, top) with StyleGaussian [26] (middle) and ARF [48] (bottom). From left to right the content resolutions are 980$\times$545 (train), 979$\times$546 (truck), and 3115$\times$2076 (counter). For the first two examples, the various outputs keep the resolution of the content, but for the HR counter scene, the output sizes are 3115$\times$2076 for SGSST, 1600$\times$1066 for StyleGaussian and 779$\times$519 for ARF (see supp. mat. for ARF results without downscaling). Thanks to its multiscale global VGG statistics, SGSST is the most faithful method regarding style consistency.
@@ 5.1 Ultra-high resolution results
Our multiscale stylization algorithm is able to transfer style details at UHR for both the content image resolution and the style image resolution. This results in unprecedentedly rich style transfer, as illustrated by Figure 3. As can be observed, minimizing the SOS loss indeed allows to decrease the style transfer loss for all scales (Figure 3 right). The approach is especially relevant when transferring the style of an UHR painting presenting style features at several scales, ranging from a specific color palette to a main curve style and to fine brushstrokes or canvas texture (see close-up views of Figure 3 and the first two lines of Figure 1). In addition, the approach is also efficient to transfer the style of a medium resolution style image to an UHR scene, as shown in the last row of Figure 1.
@@
Yet, applying style transfer at such resolutions remains computationally heavy: for the garden image of size 5187$\times$3361 (Figures 1 and 3) the style transfer takes 25.5 hours (VS 3 hours to train the initial 3DGS model [20]), that is an 8.5 overhead factor. For an image of moderate size (Figure 4 left and middle) the style transfer and the initial 3DGS training take respectively 22 minutes and 10 minutes, that is, the overhead factor is only 2.2. These time values were obtained using a single A100 GPU with 80 GB of memory. This could be accelerated by adapting the SOS loss implementation to a multi-GPU setting.
@@ 5.2 Comparison
Comparing style transfer methods is challenging because each algorithm treats the style image differently. Following high quality 2D style transfer [12, 10], our loss uses up to four scales and each scale uses five VGG layers. The UHR style images were downscaled so that the style has the same size as the content images (no upscale was applied if the style image is smaller than the content image). In contrast, ARF uses a single VGG layer from a single scale and the style image is downscaled to be twice smaller than the content input, resulting in smaller local texture. StyleGaussian reduces the style images so that is has the size 256$\times$256, a scale that is hardly sufficient to represent HR style images such as paintings.
@@
Even though there is no consensus for the quantitative evaluation of NST [17], following previous works, we report two different metrics for style transfer quality and texture consistency across views. Style transfer quality can be measured by the Gram loss [11]. A second metric proper to NeRF style transfer [25, 26, 33] is to check the short-term and long-term consistency of the radiance fields in terms of LPIPS [49] and RMSE between wrapped views. Since SGSST is the only method working with UHR images, when necessary we forcefully downgraded all the results to the resolution of ARF for comparison. The average of these two metrics over our 40 experiments is reported in Table 1. The Gram loss is the best for SGSST and, surprisingly, StyleGaussian achieves a lower Gram loss than ARF, which is not consistent with the qualitative evaluation, probably explained by ARF using use of a single VGG layer compared to five for the two other methods. In terms of consistency metrics, StyleGaussian reports to be the most stable approach followed by SGSST but note that this metric favors the lack of texture.
@@
Table 1: Quantitative comparison. Style transfer quality and texture consistency metrics averaged over 40 experiments for SGSST, StyleGaussian, and ARF. Best results are in bold, second best underlined.
@@ 5.3 Ablation study
Figure 5: Influence of optimization parameters: Allowing more 3DGS parameters to be optimized when minimizing our SOS loss does not improve the stylization quality and can dramatically impact the geometry. From left to right: Style image, Content, SGSST default (optimization of colors), results when optimizing all spherical harmonics, results when optimizing all the parameters.
@@ 6 Discussion and limitations
Our method is optimization-based and requires from several minutes to several hours of computation time depending on the resolution, as discussed in Section 5.1. On the other hand, very fast stylization for Gaussian splatting proposed by some concurrent approaches [33, 26] does not reach a satisfactory visual quality.
@@ 7 Conclusion
The supplementary material consists of the following elements:
Note that due to space constraints the images have been compressed and/or reduced in size.
Appendix A UHR style transfers of the teaser figure
Due to space limitation, style images of the main paper’s teaser figure have been displayed as tiny images regardless of their resolution. Figures 9 to 16 show the eight pairs of images of this figure in full size to better appreciate the multiscale details of the style images and their corresponding stylized results. Each style image is displayed at the same resolution as the rendered view so that one can observe that the style features are reproduced with the same size (see e.g. the stone wall of Figure 14).
@@ Appendix A UHR style transfers of the teaser figure
Appendix B Comparison experiments
@@
Figures 17 to 25 display a rendered view for each of these 40 experiments. Let us recall that for the HR scenes (Figures 17 to 21) our approach is the only one working at high-resolution. While SGSST produces outputs having the content size, StyleGaussian outputs are limited in resolution to a maximal width of 1600 or maximal height of 1200, and for ARF the content images have been downscaled by a factor 4 to obtain a low-resolution input suitable for ARF (see Section C below).
@@
Appendix C ARF and high resolution inputs

### v2

@@ 1 Introduction
In short, our approach is the first method that allows UHR style transfer directly to 3DGS. It produces high visual quality results and relies on optimizing a single multiscale loss. The simplicity of our approach ensures its reproducibility. Being optimization-based, SGSST’s main limitation is a fairly large training time that is two to eight times longer than the initial 3DGS training depending on the image resolution. Yet, considering the high quality of the results and their reproducibility, this contribution is valuable for AR/VR applications that require high visual quality for their user experience.
@@ 2 Related work
A few recent contributions show that 3DGS is a promising framework for 3D scene style transfer. Saroha et al [33] propose a solution for universal style transfer of a given 3DGS scene. The method processes the colors of Gaussians with a tiny MLP trained using fast AdaIN [15] and relying on a multi-resolution hash grid representation [29]. StyleGaussian [26] is a concurrent approach that relies on transferring encoding of VGG features to each Gaussian and applying AdaIN to these features. The new Gaussian features are then decoded into an RGB color by processing the K-nearest neighbors of the Gaussians. After training, both methods allow for instantaneous stylization with any style image, but the visual quality of the results is quite low, since high-resolution details such as brushstroke patterns are not transferred.
StylizedGS [46] is an optimization-based method that extends ARF [48] to 3DGS. It uses a training loss made of six terms combined with a color transfer preprocessing, the loss being designed to minimize changes in the 3DGS geometry while letting the style evolve via VGG NN matching. $\mathcal{G}$-Style [22] also is an optimization based approach that further uses a CLIP-based loss [31]. While these approaches produce slightly better results than ARF [48], both methods are limited in content and style resolution due to the use of nearest-neighbor matching of VGG features.
None of the current state of the art deals with HR style transfer. The stylization is either fast but very approximate due to AdaIn [33, 26] or unable to use HR style images due to NNFM [48, 46, 22]. To the best of our knowledge, SGSST is the first procedure allowing high-quality transfer for 3DGS that is trained and rendered at UHR.
@@ 3.1 Gaussian splatting representation
Starting from a multiview training set of images $\{u_{i}\}_{i=1}^{N_{\mathrm{views}}}$ of a static scene accompanied with corresponding calibrated cameras $\{\mathcal{C}_{i}\}_{i=1}^{N_{\mathrm{views}}}$ computed by structure from motion [34], the 3DGS algorithm [20] trains a set of colored 3D Gaussians $\{\mathcal{G}_{j}\}_{j=1}^{N_{\mathrm{Gaussians}}}$ so that they represent the 3D scene from any camera position. Each Gaussian $\mathcal{G}_{j}$ is represented by a finite set of features: a center position $\mu_{j}$, a covariance matrix $\Sigma_{j}$ (encoded by a scaling diagonal matrix and a rotation matrix), an opacity $\alpha_{j}$ and a view-dependent color function $c_{j}$. These parameters are used in a volumetric renderer that determines the color by summing the contribution of each Gaussian that intersects a ray with direction $(\theta,\varphi)$ via $\alpha$-blending (see e.g. [28]). The contribution of a Gaussian is its color $c_{j}(\theta,\varphi)$ multiplied by an opacity $\sigma_{j}$ defined as the maximal opacity $\alpha_{j}$ times the unnormalized Gaussian density at the ray position [20]. Thus, the resulting color $C$ for the ray is
@@ 3.2 Style transfer loss for UHR images
Figure 2: Overview of SGSST. Starting from a pretrained realistic 3DGS scene [20], we optimize the colors of each Gaussian using the new multiscale SOS loss (involving $n_{\mathrm{s}}=3$ scales in the illustration). Computing the gradient w.r.t. the loss is feasible for UHR images thanks to the SPST partition-based implementation [10]. Multiscale gradient stacking is used at the node of the rendered image to perform only one backpropagation per iteration through the 3DGS rendering pipeline.
@@
where $E_{\mathrm{content}}(x;u)=\lambda_{\mathrm{c}}\left\|V^{L_{\mathrm{c}}}(x)-V^{L_{\mathrm{c}}}(u)\right\|^{2}$, with $\lambda_{\mathrm{c}}>0$ and the style loss is defined by
@@
where $E_{\mathrm{style}}^{L}(x;v)$ is a linear combination of the mean square error between the VGG19 statistics of $V^{L}(x)$ and the one of the style features $V^{L}(v)$ [10]. While only the Gram matrices were originally used [11], it has been shown that adding control for the mean and standard deviation corrects some style transfer color artefacts [10] previously identified in the literature  [35, 32, 14].
To obtain high-quality style transfer for HR images one needs to optimize the style transfer loss using several scales and a coarse-to-fine approach [12]. Indeed, if one applies style transfer on the highest resolution only, the changes within the content image are limited to local texture and the results does not convey a painting aspect. Due to the use of VGG19 features, computing the loss $E_{\mathrm{transfer}}(x;u,v)$ and its gradient with respect to (w.r.t.) $x$ via backpropagation is memory prohibiting for UHR images. However, using a grid partition and local loss backpropagation based on precomputed global statistics, SPST [10] allows for the exact evaluation of this gradient.
@@ 4.2 Multiscale Style Transfer Loss
In the end, as illustrated by Figure 2, the stylization of UHR 3DGS is conducted by solving for
@@ 4.3 Implementation details
The 2D SPST method [10] provides the gradient of each term $E_{\mathrm{transfer}}(x^{\downarrow 2^{s}};u^{\downarrow 2^{s}},v^{\downarrow 2^{s}})$ w.r.t. to the input $x^{\downarrow 2^{s}}=\mathcal{R}(\mathcal{C}_{i};\Theta)^{\downarrow 2^{s}}$. We first backpropagate each of these gradients through the downscaling operator and stack them at the level of the rendered image $\mathcal{R}(\mathcal{C}_{i};\Theta)$. When each of the $n_{\mathrm{s}}$ scales has been treated, the stacked gradient is equal to the gradient of the full loss $E_{\mathrm{SOS}}(\mathcal{R}(\mathcal{C}_{i};\Theta);u_{i},v)$ w.r.t. to the rendered image $\mathcal{R}(\mathcal{C}_{i};\Theta)$. This gradient is then backpropagated through the Gaussian rendering pipeline to obtain the gradient w.r.t. the Gaussian colors $\Theta_{\mathrm{color}}$ (see both orange arrows in Figure 2). Using this strategy we only backpropagate one time per iteration through the 3DGS rendering pipeline instead of $n_{\mathrm{s}}$ times.
@@
$n_{\mathrm{s}}$ is set automatically to use all available scales, the coarsest resolution having sides larger than 256 for VGG19 statistics to be reliable.
@@
All our experiments have been conducted using the same SOS loss and training procedure, making our approach parameterless and fully reproducible. Our public PyTorch implementation is based on the public source codes [^1] https://github.com/graphdeco-inria/gaussian-splatting; https://github.com/bgalerne/scaling_painting_style_transfer for 3DGS [20] training and SPST [10]. Our code and videos of our results are available online [^2] Code: https://github.com/JianlingWANG2021/SGSST; Videos: https://www.idpoisson.fr/galerne/sgsst/.
@@ 5 Experiments
Figure 4: Comparison of SGSST (ours, top) with StyleGaussian [26] (middle) and ARF [48] (bottom). From left to right the content resolutions are 980$\times$545 (train), 979$\times$546 (truck), and 3115$\times$2076 (counter). For the first two examples, the various outputs keep the resolution of the content, but for the HR counter scene, the output sizes are 3115$\times$2076 for SGSST, 1600$\times$1066 for StyleGaussian and 779$\times$519 for ARF (see supp. mat. Figure 28 for ARF results without downscaling). Thanks to its multiscale global VGG statistics, SGSST is the most faithful method regarding style consistency.
@@ 5.1 Ultra-high resolution results
Our multiscale stylization algorithm is able to transfer style details at UHR for both the content image resolution and the style image resolution. This results in unprecedentedly rich style transfer, as illustrated by Figure 3. As can be observed, minimizing the SOS loss indeed allows to decrease the style transfer loss for all scales (Figure 3 right). The approach is especially relevant when transferring the style of an UHR painting presenting style features at several scales, ranging from a specific color palette to a main curve style and to fine brushstrokes or canvas texture (see close-up views of Figure 3 and the first two lines of Figure 1). To obtain such results, combining style transfer at the largest possible number of scales is critical (see ablation in supp. mat. Figure 9). In addition, the approach is also efficient to transfer the style of a medium resolution style image to an UHR scene, as shown in the last row of Figure 1.
@@
Yet, applying style transfer at such resolutions remains computationally heavy: for the garden image of size 5187$\times$3361 (Figures 1 and 3) the style transfer takes 25.5 hours (VS 3 hours to train the initial 3DGS model), that is an 8.5 overhead factor. For an image of moderate size (Figure 4 left and middle) the style transfer and the initial 3DGS training take respectively 22 and 10 minutes, that is, the overhead factor is only 2.2. These time values were obtained using a single A100 GPU with 80 GB of memory and could be accelerated by adapting the SOS loss implementation to a multi-GPU setting. Note that high computation times are inherent to UHR style transfer: Running the 2D SPST method for the 185 garden training images takes 27 hours (8.9 min. per image).
@@ 5.2 Comparison
Comparing style transfer methods is challenging because each algorithm treats the style image differently. Following high quality 2D style transfer [12, 10], our loss uses up to four scales and each scale uses five VGG layers. The UHR style images were downscaled so that the style has the same size as the content images (no upscale was applied if the style image is smaller than the content image). In contrast, ARF uses a single VGG layer from a single scale and the style image is downscaled to be twice smaller than the content input, resulting in smaller local texture. StyleGaussian reduces the style images so that it has the size 256$\times$256, a scale that is hardly sufficient to represent HR style images such as paintings.
@@
Even though there is no consensus for the quantitative evaluation of NST [17], following previous works, we report two different metrics for style transfer quality and texture consistency across views. Style transfer quality can be measured by the Gram loss [11]. A second metric proper to NeRF style transfer [25, 26, 33] is to check the short-term and long-term consistency of the radiance fields in terms of LPIPS [49] and RMSE between wrapped views. Since SGSST is the only method working with UHR images, when necessary we forcefully downgraded all the results to the resolution of ARF for comparison. The average of these two metrics over our 40 experiments is reported in Table 1. The Gram loss is the best for SGSST and, surprisingly, StyleGaussian achieves a lower Gram loss than ARF, which is not consistent with the qualitative evaluation, probably explained by ARF using a single VGG layer compared to five for the two other methods. In terms of consistency metrics, StyleGaussian reports to be the most stable approach followed by SGSST but note that this metric favors the lack of texture.
@@
Table 1: Quantitative comparison. Style transfer quality and texture consistency metrics averaged over 40 experiments for SGSST, StyleGaussian, and ARF. Best results in bold, second best underlined.
@@ 5.3 Ablation study
Figure 5: Influence of optimization parameters: Allowing more 3DGS parameters to be optimized when minimizing the SOS loss does not improve the stylization quality and can dramatically impact the geometry. From left to right: Style image, content, SGSST default (optimization of colors), results when optimizing all spherical harmonics, results when optimizing all parameters.
@@ 6 Discussion and limitations
Depending on the resolution, SGSST requires from several minutes to several hours of computation. On the other hand, fast 3DGS stylization approaches [33, 26] do not reach a satisfactory visual quality.
@@ 7 Conclusion
Acknowledgements: B. Galerne and L. Raad acknowledge the support of the project MISTIC (ANR-19-CE40-005).
@@
Our supplementary material consists of the following elements:
Note that due to space constraints all the images of this document have been compressed.
Appendix A Ablation on the number of scales
As explained in the main paper, the number of scales $n_{\mathrm{s}}$ is set automatically to use all available scales, the coarsest resolution having sides larger than 256 for VGG19 statistics to be reliable. Figure 9 presents an ablation on the number of scales $n_{s}$ showing the results for different values of $n_{s}$ and the corresponding close-ups of these results (after an initial color transfer for first 10k iterations using coarsest scale $n_{s}=4$ for all examples). One can observe that when using only the large resolution images ($n_{s}=1$) the pattern of the style transfer are limited to HR details. High-quality style transfer is only achieved when using all scales.
Figure 9: Ablation of the number of scales of the SOS loss. Style transfer results using different number of scales (starting from the same initialization obtained by 10k iterations using coarsest scale for all). High-quality style transfer is only achieved when using all scales ($n_{\mathrm{s}}=4$).
Appendix B UHR style transfers of the teaser figure
Due to space limitation, style images of the main paper’s teaser figure have been displayed as tiny images regardless of their resolution. Figures 10 to 17 show the eight pairs of images of this figure in full size to better appreciate the multiscale details of the style images and their corresponding stylized results. Each style image is displayed at the same resolution as the rendered view so that one can observe that the style features are reproduced with the same size (see e.g. the stone wall of Figure 15).
@@ Appendix B UHR style transfers of the teaser figure
Appendix C Comparison experiments
@@
Figures 18 to 26 display a rendered view for each of these 40 experiments. Let us recall that for the HR scenes (Figures 18 to 22) our approach is the only one working at high-resolution. While SGSST produces outputs having the content size, StyleGaussian outputs are limited in resolution to a maximal width of 1600 or maximal height of 1200, and for ARF the content images have been downscaled by a factor 4 to obtain a low-resolution input suitable for ARF (see Section D below). Video versions of these figures are available at: https://www.idpoisson.fr/galerne/sgsst/comparison_web.html.
@@
Moreover, we provide with Figure 27 a second version of the comparison figure (Figure 4) with close views to highlight the texture consistency of each method.
Figure 27: Comparison of SGSST (ours, top) with StyleGaussian [26] (middle) and ARF [48] (bottom) with short range views. From left to right the content resolutions are 980$\times$545 (train), 979$\times$546 (truck), and 3115$\times$2076 (counter). For the first two examples, the various outputs keep the resolution of the content, but for the HR counter scene, the output sizes are 3115$\times$2076 for SGSST, 1600$\times$1066 for StyleGaussian and 779$\times$519 for ARF (see supp. mat. for ARF results without downscaling). Thanks to its multiscale global VGG statistics, SGSST is the most faithful method regarding style consistency.
Appendix D ARF and high resolution inputs